1.separate 46 into two parts so that one part is more than twice as much as the first part. find each part.
Answer by ewatrrr(24785)


Hi
one part is more than twice as much as the first part
Let x and 2x represent the parts
Question states***
x + 2x = 46
Solving for x
3x = 46
x = 46/3. the parts are 46/3 and 92/3
CHECKING our Answer***
46/3 + 92/3 = 138/3 = 46